(3+x)/(2x-3) = -1/2
(3+x)/(2x-3) = -1/2
(3+x)/(2x-3) = -1/2
(3+x)/(2x-3) = -1/2
(3+x)/(2x-3) = -1/2
(3+x)/(2x-3) = -1/2(3+x)/(2x-3) = -1/2
 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(3+x)/(2x-3) = -1/2
2 (3 + x) = -1 (2x – 3)

6 + 2x = -2x + 3

4x = 3 -6

x = -3/4

Thanks